A 70.0-kg football player runs North at 9.0 m/s and is met by a 60.0 kg defensive tackler running South at 10.0 m/s. The total momentum is?
Answer #1

Momentum along the north, Pn = 70 * 9 kg m/s
= 630 kgm/s

Momentum along the south, Ps = 60 * 10 kg m/s (-)
= - 600 kgm/s

Total momentum = Pn + Ps
= 630 kgm/s - 600 kgm/s
= 30 kg m/s